OpenStudy (anonymous):

Find the two values of k for which y(x) = e^(kx) is a solution of the differential equation y" - 4y' + 0y = 0

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@ganeshie8

OpenStudy (chillout):

All right, first we should have \(k²*e^{kx}-4ke^{kx}=0\)

OpenStudy (chillout):

which leads to \((k-4)*k*e^{kx}=0\)

OpenStudy (chillout):

Now you are able to finish it.

OpenStudy (anonymous):

Woah woah woah, diff eq's aren't my strong suit, I'm just getting started. could you walk me into getting the k^2 * e^(kz)-4ke^(kx) = 0?

OpenStudy (anonymous):

Oops, x, not z

OpenStudy (anonymous):

Is it just doing second derivative of the function, and the first derivative, then subbing into the form given?

OpenStudy (chillout):

All right. For differential equations you must know derivatives... The differential equation in question \((y''(x)-4y'(x)+0y=0\)) is a second order linear ordinary differential equation, y''(x) is the second derivative of y(x) and y'(x) is the first derivative.

OpenStudy (chillout):

All I did was the differentiation of the given solution. Afterwards I plugged the solution in the equation.

OpenStudy (anonymous):

So with this all plugged in now, I want to solve for x = 0? Or find k in terms of x?

OpenStudy (chillout):

No, you don't want to solve in terms of x. You want find a value for k in which the left-hand side is equal to the right-hand side. (This is stated in the beggining of the question).

OpenStudy (chillout):

The question doesn't asks for values x, and most likely will never ask, unless it is a boundary value problem or initial value problem.

OpenStudy (chillout):

values of x*

OpenStudy (anonymous):

So I just want values for k that will make \[ke ^{kx}(k-4)\] equal to zero? I'm sorry if I'm still a little confused

OpenStudy (chillout):

Yes, you are right.

OpenStudy (anonymous):

0 and 4?

OpenStudy (chillout):

Yeah. Spot on!

OpenStudy (anonymous):

Wow, that was easier than I expected

OpenStudy (anonymous):

Thanks Chill!

OpenStudy (chillout):

When it gives you a solution, you will want to plug that solution and its derivatives in the ODE. That's how this kind of question is done.

OpenStudy (anonymous):

ODE being Order of Diff Eq's? Will the form given in this problem always be the same?

OpenStudy (anonymous):

Meaning y" - 4y' + 0y = 0

OpenStudy (chillout):

Nope, Ordinary Differential Equation. We have the PDEs too, but that's later for you :P

OpenStudy (anonymous):

Can I ask one more question on this topic? I've done out the work but can't see where Differential Equations fit in

OpenStudy (chillout):

Sure. Ask away.

OpenStudy (anonymous):

The solution of a certain differential equation is of the form \[y(t)=ae ^{2t} + be ^{3t}\] where a and b are constants. The solution has initial conditions y(0) = 4 and y'(0) = 3. Find the solution bu using the initial conditions to get linear equations for a and b

OpenStudy (chillout):

Show me what you have done until now.

OpenStudy (anonymous):

I did it out, and I'm pretty sure that a = -1 and b = 5. I thought I was supposed to thenn just substitute those numbers in for a and b, and call it a day, but my homework program says that's wrong

OpenStudy (anonymous):

Here's what I did

OpenStudy (anonymous):

f(0)=4, so a+b should =4, and we can get a = 4-b

OpenStudy (anonymous):

Then I found the derivative, y'(t) = 2ae^(2t)+3be^(3t)

OpenStudy (anonymous):

Subbed in 0 and set it equal to 3, where I found that 3 = 8-b, so b = 5

OpenStudy (anonymous):

Then I went back and subbed the new b into a = 4-b to get a = -1

OpenStudy (anonymous):

Submitted as an answer -1e^(2t)+5e^(3t) but webwork said it was wrong

OpenStudy (chillout):

Well, I dunno how they accept the answers, so in this one I can't help you :(

OpenStudy (chillout):

But yes, you did it right.

OpenStudy (anonymous):

Huh... Okay, well thanks again man!